OpenStudy (anonymous):

x-c/2=d solve for c ** Show Work**

jimthompson5910 (jim_thompson5910):

\[\large \frac{x-c}{2}=d\] \[\large x-c=2d\] \[\large -c=2d-x\] \[\large c=-2d+x\] \[\large c=x-2d\]